# Diffusioosmotic corner flows

### v2

@@ I Introduction
A promising approach exploits phoretic mechanisms, in which surface-generated gradients (of concentration, temperature, etc.) induce effective slip flow on confining surfaces, that in turn gives rise to bulk flow [4]. Diffusioosmosis and diffusiophoresis, the motion of particles and fluids in response to solute concentration gradients [5, 6, 7], is now a well-established active propulsion mechanism, with numerous applications in artificial active matter [8, 9].
@@ II.1 General framework
We adopt a continuum description of diffusioosmotic transport, following established theoretical frameworks [5, 6, 7], to analyse the two-dimensional flow field generated within a wedge formed by two semi-infinite lines starting from one point. The fluid within the wedge is characterised by dynamic viscosity $\eta$, density $\rho$, and contains solute of local concentration $\tilde{C}$ (number of particles per unit volume), with diffusivity $\kappa$. The chemical activity $\mathcal{A}$ of a surface $\mathcal{S}$ quantifies the fixed rate of solute release ($\mathcal{A}>0$) or absorption ($\mathcal{A}<0$) on the surface by
$\kappa\bm{n}\cdot\nabla\tilde{C}=-\mathcal{A}\quad\textrm{ on }\mathcal{S}.$ (1)
Due to the short-range interaction of solute molecules with the cavity boundary, local concentration gradients result in the motion of the solute and, consequently, drive the motion of the fluid [4]. Assuming that the thickness of the interaction layer is small compared to the dimensions of the cavity, the classical slip-velocity formulation can be used [63], and local solute gradients induce an effective slip velocity on the boundaries
@@
$\mathrm{Pe}=\frac{\mathcal{UR}}{\kappa},\qquad\text{Re}=\frac{\rho\,\mathcal{U}\mathcal{R}}{\eta}.$ (3)
@@
Given the boundary conditions that we assume in this problem, and the governing Laplace’s equation, if $\tilde{C}_{0}$ is a constant and the concentration field $\tilde{C}$ is a solution, then the excess concentration $\tilde{c}=\tilde{C}-\tilde{C_{0}}$ is a solution too. The boundary conditions (1), (2) do not depend on the absolute value of the local concentration. For all $\tilde{C}_{0}$, the resulting flow will be the same. This also implies that $\tilde{c}$ can be negative. After shifting by a positive $\tilde{C}_{0}$, we thus will still obtain a physical solution. In the next sections, we will therefore replace concentration $\tilde{C}$ with field $\tilde{c}$.
@@ III.3 Solution of Laplace and biharmonic equations in the Mellin space
The conditions for the existence of the inverse transform are that for a chosen parameter $\gamma$, the integral $\int_{0}^{\infty}\rho^{\gamma-1}c(\rho,\theta)\,d\rho$ exists and the integration line $(\gamma-i\infty,\gamma+i\infty)$ lies within the strip of analyticity of $\bar{c}(p,\theta)$. When these conditions are satisfied, the inverse is independent of $\gamma$. Unless stated otherwise, we take $\gamma=0$. Then, for appropriate $c(\rho,\theta)$, we have
@@ IV.1 A single active wall
$R(\rho)=a\rho^{\lambda}+b\rho^{-\lambda},\qquad\Theta(\theta)=p\cos\lambda\theta+q\sin\lambda\theta.$ (23)
Applying the boundary conditions, we obtain the concentration field as
@@
With this velocity profile, the problem becomes identical to the famous Taylor’s scraper [58], with the imposed slip velocity $U=MA\tan\alpha$. The stream function then has the form
@@
Figure 2: Diffusioosmotic flow induced by the activity of one phoretic wall in a wedge of angle $\alpha$ for (left to right) $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. Flow streamlines are marked in white. The colour map indicates the total velocity magnitude. The emergent bulk flow remains comparable in magnitude to the driving slip flow on tha active boundary, and decays rapidly close to the inert, no-slip wall.
@@ V.1 The case of $a\neq 0$ and $b<\infty$
$\bar{c}(p,\theta)=f_{b}-f_{a},\quad\textrm{where}\quad f_{a}(p,\theta)=A\frac{a^{p+1}}{p(p+1)}\frac{\sin(p(\alpha-\theta))}{\cos p\alpha}.$ (34)
It is assumed that $0<\alpha<2\pi$. The solution has singularities at $p^{\prime}_{k}$ such that $p^{\prime}_{k}\alpha=\frac{\pi}{2}+k\pi$ for $k\in\mathbb{Z}$, and also at $p_{0}=-1$, and a removable singularity at $p=0$.
To evaluate the concentration field, we must now invert the Mellin transform. The concentration in real space can be written as $c=I_{b}-I_{a}$, where
$I_{a}=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}A\frac{a}{p(p+1)}\frac{\sin(p(\alpha-\theta))}{\cos p\alpha}\left(\frac{a}{\rho}\right)^{p}dp.$ (35)
The integral can be evaluated by applying Cauchy’s residue theorem to a square integration contour sketched in figure 4. For $a\leq\rho$, the integration contour lies in the half-plane $\text{Real}(p)>\gamma$. For $a>\rho$, the appropriate contour lies in the $\text{Real}(p)<\gamma$ half-plane. Such a choice ensures that the integral over three sides of the square contour vanishes when the contour is extended to infinity.
@@
$p_{k}=\begin{cases}{\alpha^{-1}\left(\frac{\pi}{2}+(k-1)\pi\right)}&\text{for $k\in\mathbb{Z}_{+}$}\\ -1&\text{for $k=0$}\\ \alpha^{-1}{\left(-\frac{\pi}{2}+(k+1)\pi\right)}&\text{for $k\in\mathbb{Z}_{-}$}.\end{cases}$ (36)
@@
$I_{a}(\rho,\theta)=\begin{cases}\sum_{k\leq 0}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{for $a>\rho$}\\ -\sum_{k\geq 1}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{otherwise,}\end{cases}$ (38)
@@
Note that the residue at p=1p=-1 is responsible for satisfying the condition (33a). In the special cases of $\alpha={\pi}/{2},\pi,{3\pi}/{2}$ solutions can be obtained by taking, for example, the limit of $\alpha\to{\pi}/{2}$. The solutions for $\alpha={\pi}/{2}-\epsilon$ and $\alpha={\pi}/{2}+\epsilon$ can then be compared to check if the desired accuracy has been achieved for a chosen $\epsilon$.
@@
$c(\rho,\theta)=\begin{cases}\sum_{k\leq 0}\mathop{\mathrm{Res}}_{p=p_{k}}f_{b}&\text{for $b>\rho$},\\ -\sum_{k\geq 1}\mathop{\mathrm{Res}}_{p=p_{k}}f_{b}&\text{otherwise.}\end{cases}$ (40)
@@
In this case existence of Mellin transform requires that $\gamma\leq-1$.
We thus take $\gamma=-1-\epsilon>-{\pi}/{2}$ and write the solution as
$c(\rho,\theta)=\begin{cases}-\sum_{k\leq-1}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{for $a>\rho$},\\ \sum_{k\geq 0}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{otherwise.}\end{cases}$ (41)
@@
Figure 5: Solute concentration fields, (a)–(c), and isolines of the stream function $\psi$, (d)–(f), for an ideally absorptive wall at $\theta=\pi/6$, a phoretic wall at $\theta=0$ with a catalytic sector at $(a,b)=\{(1,3),(0,3),(1,\infty)\}$ marked in red. Here, we assumed $A=1$ and the plotted radius of the wedge is $\rho<4$. The scale bar for the absolute concentration field, $|c|$, is common for plots (a), (b) and different for panel (c).
@@ VI.1 Solution for a single catalytic sector with reflective walls
$\frac{\partial c_{0}}{\partial\theta}=-A\rho\mathds{1}_{[a_{1},b_{1}]}\quad$ $\textrm{for}\quad\theta=0$ (42a)
@@
$\bar{c}_{0}(p,\theta)=-A\frac{b_{1}^{p+1}-a_{1}^{p+1}}{p(p+1)}\frac{\cos(p(\alpha-\theta))}{\sin(p\alpha)}.$ (43)
@@
$\frac{\partial c_{\alpha}}{\partial\theta}=A\rho\mathds{1}_{[a_{1},b_{1}]}\quad$ $\textrm{for}\quad\theta=\alpha,$ (44b)
@@
$\quad\bar{c}_{\alpha}(p,\theta)=\bar{c}_{0}(p,\alpha-\theta)=-A\frac{b_{1}^{p+1}-a_{1}^{p+1}}{p(p+1)}\frac{\cos(p\theta)}{\sin(p\alpha)}.$ (45)
@@
$g_{0,a}(p,\theta)$ $=-A\frac{a}{p(p+1)}\frac{\cos(p(\alpha-\theta))}{\sin(p\alpha)}\left(\frac{a}{\rho}\right)^{p},$ (46)
$I_{0,a}$ $=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}g_{0,a}dp.$ (47)
@@
$c_{0}=I_{0,b_{1}}-I_{0,a_{1}}=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}(g_{0,b_{1}}-g_{0,a_{1}})dp.$ (48)
We now note that $g_{0,a}$ has poles at $p^{\prime}_{k}={k\pi}/{\alpha}$, with $k\in\mathbb{Z}$, and at p=1p=-1. Under the assumption that for all $k$: $p^{\prime}_{k}\neq-1$ (that is, $\alpha\neq\pi$) the residues are
@@
for $k=0:\quad\mathop{\mathrm{Res}}_{p=p_{k}}g_{a}=-\frac{Aa}{\alpha}\left[\log\left(\frac{b}{\rho}\right)-1\right],$ (49b)
@@
The choice of $\gamma$ such that $\textrm{min}\{p^{\prime}_{k}:p^{\prime}_{k}>0\}={\pi}/{\alpha}>\gamma>0$ guarantees that the Mellin transform of $c_{0}$ exists. Integrating over a square contour in one of the half-planes, we obtain
$I_{0,a}(\rho,\theta)=\begin{cases}\sum_{k\leq 0}\mathop{\mathrm{Res}}_{p=p^{\prime}_{k}}g_{a}&\text{for $a>\rho$},\\ -\sum_{k\geq 1}\mathop{\mathrm{Res}}_{p=p^{\prime}_{k}}g_{a}-\mathrm{Res}_{p=-1}g_{a}&\text{otherwise.}\end{cases}$ (50)
@@ VI.3 The diffusion problem for an arbitrary catalytic wall
$f(p,\rho,\rho^{\prime},\theta)=-\frac{\cos(p(\alpha-\theta))}{p\sin(p\alpha)}\left(\frac{\rho^{\prime}}{\rho}\right)^{p},$ (56)
@@
$\mathrm{Res}_{p=p_{k}}f=\begin{cases}-\frac{1}{\alpha}\frac{\cos(p_{k}(\alpha-\theta))}{p_{k}}\left(\frac{\rho^{\prime}}{\rho}\right)^{p}&\text{for $k\neq 0$},\\ -\frac{1}{\alpha}\log\left(\frac{\rho^{\prime}}{\rho}\right)&\text{for $k=0$}.\end{cases}$ (57)
@@
$c_{0}(\rho,\theta)=\int_{0}^{\rho}d\rho^{\prime}A(\rho^{\prime})\sum_{k\leq 0}\mathrm{Res}_{p_{k}}f(p,\rho,\rho^{\prime},\theta)-\int_{\rho}^{\infty}d\rho^{\prime}A(\rho^{\prime})\sum_{k>0}\mathrm{Res}_{p_{k}}f(p,\rho,\rho^{\prime},\theta).$ (58)

### v3

@@ I Introduction
A promising approach exploits phoretic mechanisms, in which surface-generated gradients (of concentration, temperature, etc.) induce effective slip flow on confining surfaces, which in turn gives rise to bulk flow [5]. Diffusioosmosis and diffusiophoresis, the motion of particles and fluids in response to solute concentration gradients [29, 32, 57], is now a well-established active propulsion mechanism, with numerous applications in artificial active matter [13, 63].
@@ II.1 General framework
We adopt a continuum description of diffusioosmotic transport, following established theoretical frameworks [29, 32, 57], to analyse the two-dimensional flow field generated within a wedge formed by two semi-infinite lines starting from one point. The fluid within the wedge is characterised by dynamic viscosity $\eta$, density $\rho_{0}$, and contains solute of local concentration $\tilde{C}$ (number of particles per unit volume), with diffusivity $\kappa$. The chemical activity $\mathcal{A}$ of a surface $\mathcal{S}$ quantifies the fixed rate of solute release ($\mathcal{A}>0$) or absorption ($\mathcal{A}<0$) on the surface by
$\kappa\bm{n}\cdot\nabla{\tilde{C}}=-\mathcal{A}\quad\textrm{ on }\mathcal{S},$ (1)
where $\bm{n}$ is the normal unit vector on $\mathcal{S}$. Due to the short-range interaction of solute molecules with the cavity boundary, local concentration gradients result in the motion of the solute and, consequently, drive the motion of the fluid [5]. Assuming that the thickness of the interaction layer is small compared to the dimensions of the cavity, the classical slip-velocity formulation can be used [41], and local solute gradients induce an effective slip velocity on the boundaries
@@
$\mathrm{Pe}=\frac{\mathcal{UR}}{\kappa},\qquad\text{Re}=\frac{{\rho_{0}}\,\mathcal{U}\mathcal{R}}{\eta}.$ (3)
@@
Given the boundary conditions that we assume in this problem, and the governing Laplace equation, if $\tilde{C}_{0}$ is a constant and the concentration field $\tilde{C}$ is a solution, then the excess concentration $\tilde{c}=\tilde{C}-\tilde{C_{0}}$ is a solution too. The boundary conditions (1), (2) do not depend on the absolute value of the local concentration. For all $\tilde{C}_{0}$, the resulting flow will be the same. This also implies that $\tilde{c}$ can be negative. After shifting by a positive $\tilde{C}_{0}$, we thus will still obtain a physical solution. In the next sections, we will therefore replace concentration $\tilde{C}$ with field $\tilde{c}$.
@@ III.3 Solution of Laplace and biharmonic equations in the Mellin space
The conditions for the existence of the inverse transform are that for a chosen parameter $\gamma$, the integral $\int_{0}^{\infty}\rho^{\gamma-1}c(\rho,\theta)\,d\rho$ exists and the complex integration line $(\gamma-i\infty,\gamma+i\infty)$ lies within the strip of analyticity of $\bar{c}(p,\theta)$. When these conditions are satisfied, the inverse is independent of $\gamma$. Unless stated otherwise, we take $\gamma=0$. Then, for appropriate $c(\rho,\theta)$, we have
@@ IV.1 A single active wall
$R(\rho)={r_{1}}\rho^{\lambda}+{r_{2}}\rho^{-\lambda},\qquad\Theta(\theta)={q_{1}}\cos\lambda\theta+{q_{2}}\sin\lambda\theta,$ (23)
where $r_{1}$, $r_{2}$, $q_{1}$, $q_{2}$, $\lambda$ are constants which need to be determined. Applying the boundary conditions, we obtain the concentration field as
@@
With this velocity profile, the problem becomes identical to the well-known Taylor’s scraper [72], with the imposed slip velocity $U=MA\tan\alpha$. The stream function then has the form
@@
Figure 2: Diffusioosmotic flow induced by the activity of one phoretic wall in a wedge of angle $\alpha$ for (left to right) $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. Flow streamlines are marked in white. The colour map indicates the total velocity magnitude. The emergent bulk flow remains comparable in magnitude to the driving slip flow on the active boundary, and decays rapidly close to the inert, no-slip wall.
@@ V.1 The case of $a\neq 0$ and $b<\infty$
$\bar{c}(p,\theta)=A\frac{b^{p+1}-a^{p+1}}{p(p+1)}\frac{\sin(p(\alpha-\theta))}{\cos p\alpha}.$ (34)
It is assumed that $0<\alpha<2\pi$. The solution has singularities at $p^{\prime}_{k}$ such that $p^{\prime}_{k}\alpha={\pi}/{2}+k\pi$ for $k\in\mathbb{Z}$, and also at $p_{0}=-1$, and a removable singularity at $p=0$.
To evaluate the concentration field, we must now invert the Mellin transform. The concentration in the real space can be written as $c=I_{b}-I_{a}$, where
$f_{a}$ $=A\frac{a}{p(p+1)}\frac{\sin(p(\alpha-\theta))}{\cos p\alpha}\left(\frac{a}{\rho}\right)^{p},$ (35)
$I_{a}$ $=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}f_{a}dp.$ (36)
The complex integral above can be evaluated by applying Cauchy’s residue theorem to a square integration contour sketched in figure 4. For $a\leq\rho$, the integration contour lies in the half-plane $\text{Re}(p)>\gamma$. For $a>\rho$, the appropriate contour lies in the $\text{Re}(p)<\gamma$ half-plane. Such a choice ensures that the integral over three sides of the square contour vanishes when the contour is extended to infinity.
@@
$p_{k}=\begin{cases}\displaystyle{\alpha^{-1}\left(\frac{\pi}{2}+(k-1)\pi\right)}&\text{for $k\in\mathbb{Z}_{+}$}\\ -1&\text{for $k=0$}\\ \displaystyle\alpha^{-1}{\left(-\frac{\pi}{2}+(k+1)\pi\right)}&\text{for $k\in\mathbb{Z}_{-}$}.\end{cases}$ (37)
@@
$I_{a}(\rho,\theta)=\begin{cases}\phantom{-}\sum_{k\leq 0}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{for $a>\rho$}\\ -\sum_{k\geq 1}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{otherwise,}\end{cases}$ (39)
@@
Note that the residue at p=1p=-1 is responsible for satisfying the condition (33a). In the special cases of $\alpha={\pi}/{2},\pi,{3\pi}/{2}$ solutions can be obtained by taking, for example, the limit of $\alpha\to{\pi}/{2}$. The solutions for $\alpha={\pi}/{2}-\epsilon$ and $\alpha={\pi}/{2}+\epsilon$ can then be compared to check if the desired accuracy has been achieved for a chosen small parameter $\epsilon$.
@@
$c(\rho,\theta)=\begin{cases}\phantom{-}\sum_{k\leq 0}\mathop{\mathrm{Res}}_{p=p_{k}}f_{b}&\text{for $b>\rho$},\\ -\sum_{k\geq 1}\mathop{\mathrm{Res}}_{p=p_{k}}f_{b}&\text{otherwise.}\end{cases}$ (41)
@@
In this case existence of Mellin transform requires that ${\gamma}{<}-1$.
We thus take any $\gamma$ such that $1>\gamma>-{{\pi}/{2\alpha}}$ and write the solution as
$c(\rho,\theta)=\begin{cases}-\sum_{k\leq-1}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{for $a>\rho$},\\ \phantom{-}\sum_{k\geq 0}\mathop{\mathrm{Res}}_{p=p_{k}}f_{a}&\text{otherwise.}\end{cases}$ (42)
@@
Figure 5: Solute concentration fields, (a)–(c), and isolines of the stream function $\psi$, (d)–(f), for an ideally absorptive wall at $\theta=\pi/6$, a catalytic (active) wall at $\theta=0$ with a catalytic sector at $(a,b)=\{(1,3),(0,3),({1},\infty)\}$ marked in red. Here, we assumed $A=1$ and the plotted radius of the wedge is $\rho<4$. The scale bar for the absolute concentration field, $|c|$, is common for plots (a), (b) and different for panel (c).
@@ VI.1 Solution for a single catalytic sector with reflective walls
$\frac{\partial c_{0}}{\partial\theta}=-A\rho\mathds{1}_{{[a,b]}}\quad$ $\textrm{for}\quad\theta=0$ (43a)
@@
$\bar{c}_{0}(p,\theta)=-A\frac{{b^{p+1}-a^{p+1}}}{p(p+1)}\frac{\cos(p(\alpha-\theta))}{\sin(p\alpha)}.$ (44)
@@
$\frac{\partial c_{\alpha}}{\partial\theta}=A\rho\mathds{1}_{{[a,b]}}\quad$ $\textrm{for}\quad\theta=\alpha,$ (45b)
@@
$\quad\bar{c}_{\alpha}(p,\theta)=\bar{c}_{0}(p,\alpha-\theta)=-A\frac{{b^{p+1}-a^{p+1}}}{p(p+1)}\frac{\cos(p\theta)}{\sin(p\alpha)}.$ (46)
@@
$g_{{a}}(p,\theta)$ $=-A\frac{a}{p(p+1)}\frac{\cos(p(\alpha-\theta))}{\sin(p\alpha)}\left(\frac{a}{\rho}\right)^{p},$ (47)
$I_{{a}}$ $=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}g_{{a}}dp.$ (48)
@@
$c_{0}=I_{{b}}-I_{{a}}=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}(g_{{b}}-g_{{a}})dp.$ (49)
We now note that $g_{{a}}$ has poles at $p^{\prime}_{k}={k\pi}/{\alpha}$, with $k\in\mathbb{Z}$, and at p=1p=-1. Under the assumption that for all $k$: $p^{\prime}_{k}\neq-1$ (that is, $\alpha\neq\pi$) the residues are
@@
for $k=0:\quad\mathop{\mathrm{Res}}_{p={0}}g_{a}=-\frac{Aa}{\alpha}\left[\log\left(\frac{{a}}{\rho}\right)-1\right],$ (50b)
@@
The choice of $\gamma$ such that $\textrm{max}\{p^{\prime}_{k}:p^{\prime}_{k}<0\}=-{\pi}/{\alpha}<\gamma<0$, $\gamma>-1$, guarantees that the Mellin transform of $c_{0}$ exists. Integrating over a square contour in one of the half-planes, we obtain
$I_{{a}}(\rho,\theta)=\begin{cases}\sum_{k\leq{-1}}\mathop{\mathrm{Res}}_{p=p^{\prime}_{k}}g_{a}+{\mathrm{Res}_{p=-1}g_{a}}&\text{for $a>\rho$},\\ -\sum_{k\geq{0}}\mathop{\mathrm{Res}}_{p=p^{\prime}_{k}}g_{a}&\text{otherwise.}\end{cases}$ (51)
@@ VI.3 The diffusion problem for an arbitrary catalytic wall
${h}(p,\rho,\rho^{\prime},\theta)=-\frac{\cos(p(\alpha-\theta))}{p\sin(p\alpha)}\left(\frac{\rho^{\prime}}{\rho}\right)^{p},$ (57)
@@
$\mathrm{Res}_{p=p_{k}}{h}=\begin{cases}-\frac{1}{\alpha}\frac{\cos(p_{k}(\alpha-\theta))}{p_{k}}\left(\frac{\rho^{\prime}}{\rho}\right)^{{p_{k}}}&\text{for $k\neq 0$},\\ -\frac{1}{\alpha}\log\left(\frac{\rho^{\prime}}{\rho}\right)&\text{for $k=0$}.\end{cases}$ (58)
@@
$c_{0}(\rho,\theta)=\int_{{\rho}}^{{\infty}}d\rho^{\prime}A(\rho^{\prime})\sum_{k\leq{-1}}\mathrm{Res}_{p_{k}}{h}(p,\rho,\rho^{\prime},\theta)-\int_{{0}}^{{\rho}}d\rho^{\prime}A(\rho^{\prime})\sum_{k{\geq}0}\mathrm{Res}_{p_{k}}{h}(p,\rho,\rho^{\prime},\theta).$ (59)